DETERMINATION OF THE STRUCTURE OF THE NEON METHYLENE CYCLOBUTANE COMPLEX

\underline{WEI LIN}, Department of Natural Sciences and Mathematics, University of Saint Mary, Leavenworth, KS 66048; ANDREA J. MINEI, Department of Chemistry, Wesleyan University, Middletown, CT 06459; LU KANG, Department of Natural Sciences, Union College, Barbourville, KY 40906; WALLACE C. PRINGLE and STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, CT 06459.

The pure rotational spectrum was measured and assigned for the normal and the ^{22}Ne isotopomers of the weakly bound complex, neon methylene cyclobutane. The rotational constants for the all ^{12}C isotopomer of neon methylene cyclobutane, Ne C_{5}H_{8}, are A = 3522.991(4), B = 2069.181(2), and C = 1653.733(1) MHz. The structure of Ne C_{5}H_{8} has been determined, and the coordinates of the neon in the C_{5}H_{8} principal axis systems are a = 0.272, b = 0.630, and c = 3.480 Å. The position of the neon atom is shifted 0.63 Åfrom the plane of symmetry of methylene cyclobutane due to the large amplitude motion of neon across the ring.